# Three-view Focal Length Recovery From Homographies

### v1

@@ 11.1 Calibration
To calibrate the cameras, we used a standard checkerboard pattern printed on hard plastic. We manually removed blurry or otherwise unsuitable calibration images from the dataset. We calibrated the cameras using [46]. During calibration, we used the assumption of square pixels (i.e. $f_{x}=f_{y}$). We also modeled tangential and radial distortion to obtain more accurate focal lengths. All used cameras exhibited low distortion so we use the original distorted images for evaluation to better reflect accuracy in real scenarios where cameras are expected to have low and unknown distortion. The images used for calibration, as well as the calibration code and estimated intrinsics, will be made available with the dataset.
@@ 11.2 Triplet Point Correspondence Extraction
To obtain triplet correspondences we used SuperPoint [12] with inference in the original image resolution keeping at most 2048 best keypoints. We matched the keypoints using LightGlue [34] to first perform pairwise matches. Since SuperPoint was trained to only be rotationally invariant up to $45^{\circ}$ rotations we have extracted pairwise matches by rotating one of the images four times with a step of $90^{\circ}$ and selecting the orientation which produced the largest number of matches. Afterwards, we kept only those correspondences which were matched across all three pairs thus producing triplets. We will provide the extracted correspondences as part of the dataset upon release.

### v2

@@ 11.1 Calibration
To calibrate the cameras, we used a standard checkerboard pattern printed on hard plastic. We manually removed blurry or otherwise unsuitable calibration images from the dataset. We calibrated the cameras using [zhang2000calibration]. During calibration, we used the assumption of square pixels (i.e. $f_{x}=f_{y}$). We also modeled tangential and radial distortion to obtain more accurate focal lengths. All used cameras exhibited low distortion so we use the original distorted images for evaluation to better reflect accuracy in real scenarios where cameras are expected to have low and unknown distortion. The images used for calibration, as well as the calibration code and estimated intrinsics, will be made available with the dataset.
@@ 11.2 Triplet Point Correspondence Extraction
To obtain triplet correspondences we used SuperPoint [detone2018superpoint] with inference in the original image resolution keeping at most 2048 best keypoints. We matched the keypoints using LightGlue [lindenberger2023lightglue] to first perform pairwise matches. Since SuperPoint was trained to only be rotationally invariant up to $45^{\circ}$ rotations we have extracted pairwise matches by rotating one of the images four times with a step of $90^{\circ}$ and selecting the orientation which produced the largest number of matches. Afterwards, we kept only those correspondences which were matched across all three pairs thus producing triplets. We will provide the extracted correspondences as part of the dataset upon release.